OpenStudy (anonymous):

x2 + 10x – 5 = 0 and solution is x – 5 ± √ ̅30 I understand to substitute the solution in to check but how?

OpenStudy (a_clan):

your solution should not have an 'x'

jhonyy9 (jhonyy9):

x_1,_2 =-10(+/- (100+20)*(1/2) /2 =2(-5(+/- (30)*(1/2) ) )/2 so yes after you simplified you will get your solutions - and if you like checking you need to write your equation with those solutions first with x_1 =-5+(30)*(1/2) what will be (-5+(30)*(1/2))*2 +10(-5+(30)*(1/2)) -5 =0 --- so and hence you need to get at finaly 0=0 --- and with the second solution you need doing the same -jope that you have understood it

jhonyy9 (jhonyy9):

pardon hope that you have understood it